OpenStudy (anonymous):

Question 1.1. What is the slope for the following table of values? gallons of gas used 1 2 3 4 miles driven 26 52 78 104 (Points : 1) 2 4 30 26

OpenStudy (anonymous):

2 40 30 and 26 are the answer choicces sorry

OpenStudy (texaschic101):

gallons of gas (x) -- 1 , 2 , 3 , 4 miles driven (y)---26, 52, 78, 104 now take two sets of points ....(x,y) (1,26) -- x1 = 1 and y1 = 26 (2,52) -- x2 = 2 and y2 = 52 now we use the slope formula and sub in our info slope = (y2 - y1) / (x2 - x1) slope = (52 - 26) / (2 - 1) slope = 26/1 or just 26